Hi! I need to graph this linear equation using three points: 2x+3y=8

Answer by lwsshak3(11628)

I need to graph this linear equation using three points:
use slope intercept formula: y=mx+b, m=slope=∆y/∆x, b=y-intercept
2x+3y=8
3y=-2x+8
y=-2x/3+8/3
slope=-2/3, y-intercept=8/3
start with first point, the y-intercept=8/3
from the first point, move 3 units to the right to establish the 2nd point.
from the 2nd point, move 2 units down to establish the 3rd point.
draw a straight line thru the 1st and 3rd points to graph the equation.